OpenStudy (anonymous):

A ball rolls off a horizontal table at 0.43 m/s and strikes the ground 0.20 m horizontally from the base of the table. How high is the table?

OpenStudy (anonymous):

If it hits the ground 0.2 m away, it must have taken \[0.2/0.43=0.465\] seconds to fall. For constant acceleration and starting from rest, \[s=1/2at^2\]So the distance fallen:\[s = 0.5 * (0.4651162790697674)^2 = 0.108\]Or 0.1 m

OpenStudy (anonymous):

thank you